OpenStudy (anonymous):

part A. write the expression x^2- 18 + 45 as a product of two linear expressions. Show your work and justify each step. Part B: rewrite x^2 - 6x + 9 as a square of a linear expression. Part C: do the expressions in part A and B have a common factor? PLEASE HELP ME!?

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

I am sorry, but I don't know :(

OpenStudy (anonymous):

its okay. thanks though <3 @sammixboo

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@hartnn @Preetha anybody?

OpenStudy (anonymous):

1) x^2 - 18x + 45 = 0 Find 2 numbers knowing their product is (-45) and their sum is (-b = 18). They are (-3) and (-15). Factoring form: y = (x - 3)(x - 15) 2) (x - 3) ^2 = x^2 - 6x + 9 3) Yes, (x - 3)

OpenStudy (anonymous):

Sorry, correction: 2 numbers knowing their product is (45)

OpenStudy (anonymous):

@thu1935 thanks so much!!!